Stock A is expected to return 12 percent in a normal economy and lose 7 percent in a recession. Stock B is expected to return 8 percent in a normal economy and 2 percent in a recession. The probability of the economy being normal is 80 percent and the probability of a recession is 20 percent. What is the covariance of these two securities?

Group of answer choices

.003876

.004203

.003280

.004115
Answer #1

Average return for Stock A:
(0.80*1.12) + (0.20*0.93)
=0.896+0.186
=1.082

Average return for stock B:
(0.80*1.08) + (0.20*1.02)
=0.864+0.204
=1.068

So now calculating co-variance
=Sum(Xi-Xm)(Yi-Ym)/n-1
=Sum((0.80*1.12-1.082)(0.80*1.08-1.068)(0.20*0.93-1.082)(0.20*1.02-1.068)/2-1
=0.003876